# Superposition in Graph Neural Networks

### v1

@@ Appendix A Generalized Mean Pooling
Let $N$ be the number of nodes, $x_{id}\in\mathbb{R}$ the $d$-th dimension of node $i$, and $p\in\mathbb{R}$ a pooling parameter. Introduce a small stabiliser $\varepsilon>0$ and define
@@
1. Element-wise transformation: For every node $i$ and feature dimension $d$,
@@ 2. Pooling: Sum the transformed features and take their mean,
3. Inverse transformation: Apply the signed $p$-th root to obtain the pooled feature
@@
The resulting graph-level representation is the vector
@@
Appendix B Symmetry Breaking
From Appendix A we note that the pooled elements are of the form:
@@ Appendix B Symmetry Breaking
Applying the chain rule gives:
$\frac{\partial h_{d}}{\partial x_{id}}=\frac{1}{N^{1/p}}\lvert s_{d}\rvert^{\frac{1}{p}-1}\operatorname{sgn}(s_{d})\times\lvert x_{id}\rvert^{p-1}\operatorname{sgn}(x_{id}).$
Appendix C SVD-Analysis
Let $H\in\mathbb{R}^{n\times d}$ be the matrix whose columns collect the averaged $d$-dimensional pooled embeddings of single-active-feature graphs, where $n$ is the number of features. Performing a singular–value decomposition,
$H\;=\;U\,\Sigma\,V^{\top},$
we obtain
  • $U\in\mathbb{R}^{n\times n}$: an orthogonal matrix whose columns $u_{1},\dots,u_{n}$ are the left singular vectors $\bigl{(}U^{\top}U=I_{n}\bigr{)}$;
  • $\Sigma\in\mathbb{R}^{n\times d}$: a (generally rectangular) diagonal matrix
    $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0),\qquad\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0,$
    where $r=\operatorname{rank}(H)$ and the positive entries $\sigma_{i}$ are the singular values;
  • $V\in\mathbb{R}^{d\times d}$: an orthogonal matrix whose columns $v_{1},\dots,v_{d}$ are the right singular vectors $\bigl{(}V^{\top}V=I_{d}\bigr{)}$.
Interpretation for pooled embeddings.
In our context, the directions $v_{i}$ in ascending order capture the most dominant linear subspaces along which the pooled embeddings vary when a single graph feature is activated. Hence, if any $\sigma_{i}=0$, the embeddings lie in a lower dimensional subspace.
Appendix D Sharp SI transitions during transition
A training-time phase transition.
On PAIRWISE we observe two training modes: (i) full-rank runs where $r_{\tau}(H)$ rises to $d$ and $\mathrm{EffRank}(H)$ tracks it; (ii) low-rank runs where $r_{\tau}(H)$ remains <d<d for most of training. Across seeds, SI/WNO of the centroids exhibit a reproducible four-stage pattern:
@@ Appendix D Sharp SI transitions during transition
Appendix E Geometric Regimes for Anti-Aligned Embedding Directions
E.1 Problem Formulation
We first consider the case where the feature number $n$ is small and no hyperplane has formed: Let $\{\,\mathbf{v}_{i}\in\mathbb{R}^{d}\mid\|\mathbf{v}_{i}\|_{2}=1,\;i=1,\dots,n\}$ be the (column-normalized) embedding directions learned by the network’s last hidden layer, and let $\{\mathbf{w}_{i}\}_{i=1}^{n}$ denote the corresponding unit-length rows of the classifier weight matrix. The anti-alignment hypothesis we explored in Section 6 states that after training
@@ E.1 Problem Formulation
i.e. every embedding is (almost) identical to its class weight, and pairwise inner products are non-positive if possible. In other words, the network attempts to realise a set of $n$ mutually obtuse vectors in $\mathbb{R}^{d}$. The feasibility of (A.1) depends on the relation between $n$ and $d$, which can be separated into four distinct regimes.
E.2 Regime Analysis
@@
    Over-complete case, $n>2d$. Via a simple counting argument, at least two vectors must lie in the same closed half-space; hence their inner product is >0>0, so one pair is necessarily acute.
@@
    Simplex threshold, $n=d+1$. The largest set of vectors that can satisfy $\mathbf{v}_{i}^{\!\top}\mathbf{v}_{j}<0$ for all $i\neq j$ is the vertex set of a centred regular $d$-simplex. This is the first value of $n$ that permits pairwise obtuse directions.
@@
    Under-complete case, $n\leq d$. One can choose any $n$ orthogonal unit vectors (or the $n$ vertices of a degenerate $n$-simplex in a n1n-1 dimensional subspace).
E.3 Implication for Rank Collapse
Assume training has converged to a low-rank subspace $\mathcal{S}\subset\mathbb{R}^{d}$ with $\dim(\mathcal{S})=r<d$ and all $\mathbf{v}_{i}\in\mathcal{S}$. Activating a new latent dimension $\mathbf{e}_{\perp}\perp\mathcal{S}$ perturbs one vector as $\mathbf{v}_{k}\mapsto\tilde{\mathbf{v}}_{k}=\sqrt{1-\varepsilon^{2}}\,\mathbf{v}_{k}+\varepsilon\mathbf{e}_{\perp}$. Then, for $i\neq k$,
@@ E.3 Implication for Rank Collapse
because $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}\leq 0$ in the metastable state. Hence the angle decreases [^4] Equality holds only when $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}=0$, keeping the angle unchanged at $90^{\circ}$. towards $90^{\circ}$, which reduces pairwise obtuseness. As discussed in Section 6 this is is penalised by the BCE loss function. During back propagation gradients oppose the escape from the low-rank basin, effectively creating a metastable minima. If $\varepsilon$ grows large enough performance gains may begin to outweigh the angle penalty, and the rank grows.
E.4 Rank Collapse with Hyperplane
@@
yet when a separating hyperplane emerges, the learned embeddings $\{\mathbf{v}_{i}\}$ and the classifier weights $\{\mathbf{w}_{i}\}$ are no longer almost identical. Empirically, $\mathbf{v}_{i}$ are often acute rather than mutually obtuse, but the cross-class dot products stay strictly negative:
@@
since $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{k}<0$ by (A.4) while $\mathbf{w}_{j}^{\!\top}\mathbf{e}_{\perp}=0$. Thus the cross-class dot product becomes less negative, which increases BCE loss. In short, even after the hyperplane forms, the network remains trapped in a low-rank metastable basin for the same reason as before: cross-class obtuseness ($\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$) resists escapes that would reduce the margin.
E.5 Singular-Value Thresholds
Here we make a few general remarks about the threshold for singular values, which signals the end of the metastable regime.
  1. 1.
    We expect from our theoretical description in Appendix E.3, that the threshold value depends on both the number of active features and the rank of the associated representation. This is because the more obtuse active features are, the stronger the effect that creates the energy barrier.
  2. 2.
    We empirically find the threshold to be approximately $10^{-4}$, based on a sample size of 50.
  3. 3.
    The threshold characteristics change noticeably with the active feature number and the rank of the associated representation.
E.6 Leaky ReLU Rank

### v2

@@ Appendix C Generalized Mean Pooling
Let $N$ be the number of nodes, $x_{id}\in\mathbb{R}$ the dd‑th dimension of node $i$, and $p\in\mathbb{R}$ a pooling parameter. Introduce a small stabiliser $\varepsilon>0$ and define
@@
1. Element‑wise transformation: For every node $i$ and feature dimension $d$,
@@ 2. Pooling: Sum the transformed features and take their mean,
3. Inverse transformation: Apply the signed pp‑th root to obtain the pooled feature
@@
The resulting graph‑level representation is the vector
@@
Appendix D Symmetry Breaking
Recall the stabilized generalized mean pooling from App. C. For each feature dimension $d$, let $x_{id}\in\mathbb{R}$ be the feature of node $i$, and define
$\tilde{x}_{id}=g_{p,\varepsilon}(x_{id})=\operatorname{sgn}(x_{id})\bigl(|x_{id}|+\varepsilon\bigr)^{p},\qquad s_{d}=\sum_{i=1}^{N}\tilde{x}_{id},\qquad\bar{s}_{d}=\frac{1}{N}s_{d}.$
The pooled feature is then
@@ Appendix D Symmetry Breaking
We now compute $\partial h_{d}/\partial x_{id}$, away from the measure-zero points where $x_{id}=0$ or $\bar{s}_{d}=0$ (so that the sign functions are locally constant).
Step 1: derivative of the inner transform.
For fixed $d$ and node $i$, write
$\tilde{x}_{id}=\operatorname{sgn}(x_{id})\bigl(|x_{id}|+\varepsilon\bigr)^{p}.$
Using $\frac{\partial|x_{id}|}{\partial x_{id}}=\operatorname{sgn}(x_{id})$ and $\operatorname{sgn}(x_{id})^{2}=1$, we obtain
$\frac{\partial\tilde{x}_{id}}{\partial x_{id}}=\operatorname{sgn}(x_{id})\cdot p\bigl(|x_{id}|+\varepsilon\bigr)^{p-1}\cdot\frac{\partial|x_{id}|}{\partial x_{id}}=p\bigl(|x_{id}|+\varepsilon\bigr)^{p-1}.$
Hence
$\frac{\partial s_{d}}{\partial x_{id}}=\frac{\partial}{\partial x_{id}}\Bigl(\sum_{j=1}^{N}\tilde{x}_{jd}\Bigr)=\frac{\partial\tilde{x}_{id}}{\partial x_{id}}=p\bigl(|x_{id}|+\varepsilon\bigr)^{p-1},$
and therefore
$\frac{\partial\bar{s}_{d}}{\partial x_{id}}=\frac{1}{N}\frac{\partial s_{d}}{\partial x_{id}}=\frac{p}{N}\bigl(|x_{id}|+\varepsilon\bigr)^{p-1}.$
Step 2: derivative of the outer map.
Define $y=\bar{s}_{d}$ and write
$h_{d}=f(y)=\operatorname{sgn}(y)\bigl(|y|+\varepsilon\bigr)^{1/p}.$
For $y\neq 0$, $\operatorname{sgn}(y)$ is locally constant, and we have
$\frac{\partial h_{d}}{\partial y}=\operatorname{sgn}(y)\cdot\frac{1}{p}\bigl(|y|+\varepsilon\bigr)^{1/p-1}\cdot\frac{\partial|y|}{\partial y}=\frac{1}{p}\bigl(|y|+\varepsilon\bigr)^{1/p-1},$
since $\frac{\partial|y|}{\partial y}=\operatorname{sgn}(y)$ and the sign factors cancel.
Step 3: chain rule.
Combining the two steps by the chain rule,
$\frac{\partial h_{d}}{\partial x_{id}}=\frac{\partial h_{d}}{\partial\bar{s}_{d}}\cdot\frac{\partial\bar{s}_{d}}{\partial x_{id}}=\frac{1}{p}\bigl(|\bar{s}_{d}|+\varepsilon\bigr)^{1/p-1}\cdot\frac{p}{N}\bigl(|x_{id}|+\varepsilon\bigr)^{p-1}.$
Thus the stabilized gradient of generalized mean pooling is
$\boxed{\frac{\partial h_{d}}{\partial x_{id}}=\frac{1}{N}\bigl(|\bar{s}_{d}|+\varepsilon\bigr)^{\frac{1}{p}-1}\bigl(|x_{id}|+\varepsilon\bigr)^{p-1}}$
for all $x_{id}$ and $\bar{s}_{d}$ away from the non-differentiable points of the absolute-value function.
Remark. If we further let $\varepsilon\to 0$ and assume non-negative activations so that $\bar{s}_{d}=s_{d}/N$ and $x_{id}\geq 0$, this reduces to $\frac{\partial h_{d}}{\partial x_{id}}=N^{-1/p}|s_{d}|^{1/p-1}x_{id}^{p-1}$.
Appendix E Sharp SI transitions during transition
A training‑time phase transition.
On PAIRWISE we observe two training modes: (i) full‑rank runs where $r_{\tau}(H)$ rises to $d$ and $\mathrm{EffRank}(H)$ tracks it; (ii) low‑rank runs where $r_{\tau}(H)$ remains <d<d for most of training. Across seeds, SI/WNO of the centroids exhibit a reproducible four‑stage pattern:
@@ Appendix E Sharp SI transitions during transition
Appendix F Geometric Regimes for Anti‑Aligned Embedding Directions
F.1 Problem Formulation
We first consider the case where the feature number $n$ is small and no hyperplane has formed: Let $\{\,\mathbf{v}_{i}\in\mathbb{R}^{d}\mid\|\mathbf{v}_{i}\|_{2}=1,\;i=1,\dots,n\}$ be the (column‑normalized) embedding directions learned by the network’s last hidden layer, and let $\{\mathbf{w}_{i}\}_{i=1}^{n}$ denote the corresponding unit‑length rows of the classifier weight matrix. The anti‑alignment hypothesis we explored in Section 6 states that after training
@@ F.1 Problem Formulation
i.e. every embedding is (almost) identical to its class weight, and pairwise inner products are non‑positive if possible. In other words, the network attempts to realise a set of $n$ mutually obtuse vectors in $\mathbb{R}^{d}$. The feasibility of (A.1) depends on the relation between $n$ and $d$, which can be separated into four distinct regimes.
F.2 Regime Analysis
@@
    Over‑complete case, $n>2d$. Via a simple counting argument, at least two vectors must lie in the same closed half‑space; hence their inner product is >0>0, so one pair is necessarily acute.
@@
    Simplex threshold, $n=d+1$. The largest set of vectors that can satisfy $\mathbf{v}_{i}^{\!\top}\mathbf{v}_{j}<0$ for all $i\neq j$ is the vertex set of a centred regular dd‑simplex. This is the first value of $n$ that permits pairwise obtuse directions.
@@
    Under‑complete case, $n\leq d$. One can choose any $n$ orthogonal unit vectors (or the $n$ vertices of a degenerate nn‑simplex in a n1n-1 dimensional subspace).
F.3 Implication for Rank Collapse
Assume training has converged to a low‑rank subspace $\mathcal{S}\subset\mathbb{R}^{d}$ with $\dim(\mathcal{S})=r<d$ and all $\mathbf{v}_{i}\in\mathcal{S}$. Activating a new latent dimension $\mathbf{e}_{\perp}\perp\mathcal{S}$ perturbs one vector as $\mathbf{v}_{k}\mapsto\tilde{\mathbf{v}}_{k}=\sqrt{1-\varepsilon^{2}}\,\mathbf{v}_{k}+\varepsilon\mathbf{e}_{\perp}$. Then, for $i\neq k$,
@@ F.3 Implication for Rank Collapse
because $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}\leq 0$ in the metastable state. Hence the angle decreases [^4] Equality holds only when $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}=0$, keeping the angle unchanged at $90^{\circ}$. towards $90^{\circ}$, which reduces pairwise obtuseness. As discussed in Section 6 this is is penalised by the BCE loss function. During back propagation gradients oppose the escape from the low‑rank basin, effectively creating a metastable minima. If $\varepsilon$ grows large enough performance gains may begin to outweigh the angle penalty, and the rank grows.
F.4 Rank Collapse with Hyperplane
@@
yet when a separating hyperplane emerges, the learned embeddings $\{\mathbf{v}_{i}\}$ and the classifier weights $\{\mathbf{w}_{i}\}$ are no longer almost identical. Empirically, $\mathbf{v}_{i}$ are often acute rather than mutually obtuse, but the cross‑class dot products stay strictly negative:
@@
since $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{k}<0$ by (A.4) while $\mathbf{w}_{j}^{\!\top}\mathbf{e}_{\perp}=0$. Thus the cross‑class dot product becomes less negative, which increases BCE loss. In short, even after the hyperplane forms, the network remains trapped in a low‑rank metastable basin for the same reason as before: cross‑class obtuseness ($\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$) resists escapes that would reduce the margin.
F.5 Leaky ReLU Rank